# Tracking electron capture processes in classical molecular dynamics simulations for spectral line broadening in plasmas

### v1

@@ Abstract
Plasma spectroscopy is a fundamental tool for diagnosing laboratory and astrophysical plasmas. Accurate interpretation of spectra depends upon precise modeling and comprehension of Stark-broadening and other mechanisms affecting spectral lines. In this context, computer simulations have emerged as invaluable tools, offering idealized experiments with well-defined conditions. Molecular dynamics simulations, in particular, excel at replicating the particle interactions within the plasma and their impact on the state of a radiating atom or ion. However, these simulations present challenges in tracking electron capture processes, since setting an unambiguous criterion to distinguish between bound and free electrons is not trivial. In this paper we introduce a new algorithm that, within a classical framework, precisely identifies the scenario in which an electron is captured by an ion and then follows a stable orbit around it. The algorithm’s applicability extends to emitters with charges $Z\geq 1$. Importantly, the procedure enables the correct identification of valid time-histories of the electric microfield perturbing the emitting ion, which will be used for subsequent line shape calculations. As a result, our simulations naturally and accurately incorporate the effects of both strong collisions and electron capture phenomena on spectral line broadening.
@@ I Introduction
Accurate modeling of spectral line shapes is indispensable for investigating the properties of high energy density plasmas. The analysis of emission or absorption spectra, which typically involves the comparison of experimental data with spectroscopic models, stands as a key diagnosis tool for laboratory and astrophysical plasmas. Spectral line shapes, broadened and shifted by the kinetics of plasma particles and their interactions, provide relevant information about the plasma temperature and density; and play a pivotal role in opacity calculations, crucial for radiation hydrodynamics simulations.
@@
Spectral lines arise from atomic transitions between bound energy states of the radiator. In the context of hot and dense plasmas the Stark effect emerges as the dominant broadening mechanism [1]. Unlike the isolated scenario, when embedded in a plasma, the radiator is subject to the electric field generated by its surrounding charged particles during the radiation time interval. The shape of the spectral lines in this case depends on the plasma-induced electric field, its time evolution (tied to the the plasma particle dynamics) and the radiators’ ensemble.
@@
Despite significant strides in recent decades [2, 3], modeling spectral line shapes remains a formidable challenge owing to its interdisciplinary nature, requiring the integration of various ingredients from atomic physics, plasma physics, statistical mechanics, collision theory, and computational physics. Moreover, in the high-density regime, dedicated validating experiments are notably scarce, presenting a substantial obstacle to grounding theoretical models in empirical reality. In response to these challenges, the Spectral Line Shapes in Plasmas workshop series [4, 5, 6, 7] has emerged as a valuable initiative, providing a platform to compare the multiple approaches to the problem, identify sources of discrepancies, and scrutinize the validity of different approximations. Such collaborative endeavors have stimulated the specialized community to undertake a deeper exploration of missing physics in line broadening theory and to improve the reliability of the existing models.
@@
Two primary approaches are used for line-broadening calculations: analytic methods and computer simulations. Analytic models navigate the complexities of line broadening by incorporating assumptions and mathematical manipulations to handle the stochastic nature of line emission. One common approximation involves the separate calculation of electron and ion broadening contributions. Since ions are much heavier than electrons, they are assumed to be static over the relevant electron time scales. The analytic approach then puts the effort in obtaining an accurate representation of the electron broadening operator. In this regard, some common approximations include the simple and well-known impact approximation, or the second-order approximation as a simplified way to compute the electron broadening operator. Once the electron broadening contribution is computed for a given ion distribution, i.e. for a given value of the ion microfield, the line emission is integrated over the statistical distribution of the ion microfield for the conditions of interest, a facet that entails its own extensive research area in microfield probability distributions. The inherent advantages of analytic methods lie in their computational efficiency, as the time required to calculate a line shape is generally much shorter than for computer simulations, and they operate without numerical noise.
On the other hand, simulations prove particularly advantageous in gaining insights into the line broadening problem, excelling at replicating plasma particle interactions and their impact on the state of a radiating atom or ion. Simulated plasma particles traverse within a confined space, following classical trajectories. These classical particles generate a time-dependent Coulomb potential or, equivalently, a time-sequence of the electric field, which is then utilized to solve for the time evolution of the radiator, subsequently generating a spectrum. Simulations often simplify the problem and reduce computational time by employing straight-line paths or hyperbolic trajectories for particle movement. Noteworthy exceptions to this norm are molecular dynamics simulations, such as the ones presented in this work, which uniquely account for full N-body interactions.
It is in this context that our work is framed. We present an algorithm to accurately track the electron capture processes within molecular dynamics computer simulations, key to obtain the correct time-sequences of the electric field within the plasma and accurately account for the additional recombination broadening.
This manuscript is structured as follows: in Section II, we give an overview of how computer simulations in the context of line broadening calculations work and present the need for an algorithm capable of detecting electron capture process. Section III details the criterion itself. An example of the application of this criterion to a computer simulation is presented in Sec. IV.1, and a comparison of our results with an additional ionization metric is given in Sec. IV.2. Finally, Sec. V summarizes the conclusions and limitations of this work.
II Overview of computer simulations
Computer simulations for calculating Stark-affected spectra usually follow these steps [8]:
  • a)
    In a computer, one reproduces a plasma composed of charged and/or neutral particles which are distributed in a given region of space. These particles move according to certain pre-established statistical distributions which arise from a given physical configuration (electron density and temperature). The simulation evolves in discrete timesteps and the calculation permits the description of the position and velocity of the particles at every timestep.
  • b)
    The simulated plasma includes the emitters that are considered for the spectral calculations. For each timestep, the electric field produced by all plasma particles on a given emitter is calculated. This is the time-dependent field sequence that this work focuses on.
  • c)
@@ II Overview of computer simulations
    The emitting atom or ion internally evolves, subjected to the effect of this electric field. By numerically solving the time-dependent Schrödinger equation, one obtains the corresponding evolution operator. It is then possible to calculate the dipolar moment of said emitter as a function of time. The Fourier transform of the autocorrelation function of this dipolar moment corresponds to the emission spectrum from the emitter under consideration.
  • d)
    This process is repeated a sufficient number of times. The average value of the individually produced spectra corresponds to the spectrum of interest.
Let us now consider exclusively the first step of the process described above: the numerical treatment of the simulated plasma. Generally, we can consider two types of computer simulations: those where the particles are considered physically and statistically independent, and those which take into account the interactions among charged particles. In the first case, the idea is to numerically reproduce a plasma of free and independent particles that move within the plasma following rectilinear trajectories with fixed velocities. We will refer to these methods as Trivial Molecular Dynamics (TMD). In the second case, the calculation takes into account the electric interaction of all particles in the plasma. It is a more accurate method (and more computationally expensive) which must solve Newton’s equations of motion to determine the trajectories of all particles. We will refer to these methods as Full Molecular Dynamics (FMD).
TMD calculations, much more computationally affordable, are useful when the typical electric-interaction energies are much lower than the mean kinetic energy of particles (a weakly-coupled high temperature plasma) and, specially for non-charged emitters [9, 10, 11, 12, 13, 14]. The second method (FMD) is necessary for medium and strongly-coupled plasmas and, specially when the emitter is an ion, given that its interaction with the surrounding particles in the plasma affects both its own trajectory and those of the other particles, thus modifying the time-evolution of the electric field of interest [15, 16, 17]. In this work, we focus on this type of calculations.
@@
It is worth mentioning that there is an intermediate type of calculations, in which the perturbers are treated as independent particles, but their interaction with the emitter is taken into account explicitly. This results in the charged perturbers moving in hyperbolic trajectories when a Coulomb field is considered, or in slightly different paths if Debye screening is introduced, in which case the corresponding differential equation must be solved numerically [18].
In a FMD simulation, the Coulombian interaction between the plasma particles is considered clasically. To avoid divergencies in the electron-ion interaction energies at short distances, it is necessary to regularize the interaction potential. Several procedures for this regularization have been reported (see Bonitz et al. [19] and references therein), and in this work we will use that described in Gigosos et al. [20], which considers a parabolic potential for interparticle distances below a given value $a$ and hyperbolic for larger distances. This procedure naturally accounts for the diffraction of electrons around an ion and determines an effective atom size.
@@
During the evolution of the simulation, it is possible that the binding energy (negative) between an electron and an ion becomes too large - in absolute value - with respect to their kinetic energies, so that the electron becomes trapped by the ion. In that case, the electron becomes bound and its trajectory follows an orbit around the ion. The two particles in this pair can therefore no longer be considered free and must be labeled as a neutral atom, or as an ion with a positive charge lower than that of the free ion. This circumstance has been extensively discussed in our previous work [20], where we developed a statistical model for the equilibrium of the simulated plasma, taking into account the process of electron capture and release by the ions (recombination and ionization respectively). The model perfectly reproduces what is observed in the simulation, therefore allowing us to correctly determine the density of free electrons in the plasma, considering that these trapped electrons do not contribute to its value. For this model, it was necessary to establish a clear criterion to determine when an electron-ion pair must be considered bound, and when those particles must be considered free. While in Gigosos et al. [20] ions of charge $+q_{e}$ were used (where $q_{e}$ is the electron charge), in this work we are interested in plasmas with charge $+Zq_{e}$, with $Z>1$. The purpose of this work is then to extend this criterion to situations in which the number of trapped electrons can exceed $1$.
@@
Furthermore, there is another important reason that requires knowing when an ion has captured an electron and when it can be considered free. Let us consider a FMD simulation in which we are interested in calculating the spectrum from an ionic emitter with charge $+Zq_{e}$. To that end, during the simulation process, we must register the temporal sequence of the electric field seen by that charged emitter, which is produced by the ions and free electrons surrounding it. If, at a given moment, an electron is captured by the ion of consideration, from that moment on, the electron will orbit around it at a short distance. Therefore, the electric field felt by that emitter will be an oscillating function with a large amplitude (since the oscillating electron is very close), which spoils the field sequence. In addition, this ion, originally with charge $+Zq_{e}$ has now become, along with the trapped electron, an ion of charge $+(Z-1)q_{e}$ and can therefore no longer be considered a contributor to the spectral line in which we are interested. The moment the electron is captured determines when the ion changes its species and therefore completely breaks the coherence of the spectral line that was being emitted. This problem was already considered in Gigosos et al. [17], where the process is discussed but no details are given on the criterion that the computer follows to determine when an electron can be considered bound or truly free. This is the purpose of this work.
III Criterion for detecting trapped electrons in a computer simulation
As mentioned before, detecting ionization and recombination processes is of fundamental importance in computer simulations. In these calculations, when the ion whose spectrum is being calculated captures an electron, it no longer belongs to the ion species of interest and, therefore, its internal evolution is not relevant: the emission coherence is abruptly lost. This causes a broadening of the line by the recombination process, which highlights one the advantages of these calculations [17]. As a consequence, in practice, the corresponding time-sequence of the electric field is cut in the simulation process. It is necessary then to wait for the corresponding ion to recover its original charge in order to continue using it for the calculation.
@@ III Criterion for detecting trapped electrons in a computer simulation
To correctly select the valid field sequences we choose a global criterion, in which whether an ion has trapped an electron or not, not only depends on its conditions at a given time, but also on its history. For a given ion, we store (for each step of the simulation) its potential energy, the electric field at its position and the label (an index identifying each electron in the simulation) of the $Z+1$ closest electrons, where $Zq_{e}$ is the charge of the emitter. For this work, we rescaled the ion-electron potential energy defined in Gigosos et al. [20] for $Z=1$ ions, to
$V_{ie}(r)=\begin{cases}V_{i}\left[\frac{1}{3}\left(\frac{r}{a}\right)^{2}-1\right],&\quad r\leq a\\ -\frac{Zq_{e}^{2}}{4\pi\varepsilon_{0}}\frac{1}{r},&\quad a<r\leq R_{I}\\ 0,&\quad r>R_{I}\end{cases}$ (1)
where $V_{i}$ is the ionization energy corresponding to the ion of charge $Z$ and $R_{I}$ is the radius of the interaction sphere, defined as half the size of the cubic simulation box. For ions with $Z>1$, $a$ is also rescaled as
$a=\frac{3}{2}\frac{Zq_{e}^{2}}{4\pi\varepsilon}\frac{1}{V_{i}}.$ (2)
The first step then is to track down the moments when the potential energy of the emitter is lower than a given value. In particular, we chose a threshold energy of
$\mathscr{E}_{th}=-\frac{3}{4}V_{i}.$ (3)
Whenever the ion energy is lower than $\mathscr{E}_{th}$, the electron that is closest to it (its first neighbor) is identified. Then, the time interval during which that electron has been among the $Z+1$ closest neighbors is considered. If this interval is longer than a given threshold value, then the average total energy of the electron-ion pair in that interval is calculated. If the pair has a negative total energy during that interval, then the electron is considered to be bound for the whole interval. If not, it is considered that the negative energy seen by the emitter is caused by a strong collision, since the electron causing it did not remain around the ion. It is worth noting that the $Z+1$ closest neighbors are considered, since it is possible that a fast electron gets closer to the ion core than other bound electrons for a short period of time, without freeing them.
Since the full time-history of the ion is needed in order to check whether an electron is bound or free, its history must be saved to files in the disk. In this way, it can be accessed when the energy criterion is met, and the correct field sequences can be extracted.
Determining the timescale of bound electrons in the simulation
@@
$\tau_{T}=\frac{2\pi r_{T}}{v_{T}}=\frac{Zq_{e}^{2}}{2\varepsilon_{0}}\cdot\frac{m_{e}^{1/2}}{(2k_{B}T_{e})^{3/2}},$ (6)
or, in simulation units111 The units that the simulation works with are important. The relation between the numbers in the simulation and physical magnitudes is determined by two parameters in the simulation: (1) the coupling degree $\rho$, which is a non-dimensional parameter whose value is fixed prior to the simulation and corresponds to the relation between the typical distance between electrons and the Debye radius; and (2) the value of the ionization potential in simulation units (which is also fixed). The unit of length is taken as the typical distance between electrons (free or not), the unit of time is the characteristic time in the electron dynamics, and the unit of mass is that of the electron. In this manuscript, we work with those units. $\mathscr{E}_{0}$ is the corresponding unit of energy, and $E_{0}$ that of the electric field. For a more detailed discussion on the units in the simulation, see Gigosos et al. [20].
$\tau_{T}^{S.U.}=\frac{\pi}{3}Z\rho^{2}$ (7)
We take this characteristic orbit time as a basis to consider whether an electron is bound to an ion. Although one might be tempted to consider $\tau_{T}$ as the temporal threshold - that is, if an electron is within the $Z+1$ closest neighbors of a given ion for at least $\tau_{T}$, and its total energy is negative, it is considered bound - this criterion would overestimate the number of bound electrons, since an electron can loop around an ion once without being bound. For this reason, we choose $\tau_{th}=3\tau_{T}$ as the threshold for considering an electron as bound to a given ion.
We note here that the above expression for $\tau_{T}$ was obtained for a hyperbolic potential. However, as indicated in Eqn. 1, in our simulations, the potential becomes parabolic for distances $r<a$, and thus, care must be taken to ensure that $r_{T}>a$. By comparing the above expressions for both $r_{T}$ and $a$, it can easily be seen that this condition is equivalent to
@@
Figure 1: Time-history of a plasma with $\rho=0.6$ and emitters with charge 2. From top to bottom: potential energy of an emitter as a function of time, indices of the three closest electrons at each moment, and electric field. The shaded regions correspond to the moments when a particle is trapped (see footnote 1).
IV Results and discussion
IV.1 Application of the criterion

### v2

# Tracking electron capture process in classical molecular dynamics simulations for spectral line broadening in plasmas

@@ Abstract
Plasma spectroscopy is a fundamental tool for diagnosing laboratory and astrophysical plasmas. Accurate interpretation of spectra depends upon precise modeling and comprehension of Stark-broadening and other mechanisms affecting spectral lines. In this context, computer simulations have emerged as valuable tools, offering idealized experiments with well-defined conditions. Molecular dynamics simulations, in particular, excel at replicating the particle interactions within the plasma and their impact on the state of a radiating atom or ion. However, these simulations present challenges in tracking electron capture processes, since setting an unambiguous criterion to distinguish between bound and free electrons is not trivial. In this paper we introduce a new algorithm that, within a classical framework, precisely identifies the scenario in which an electron is captured by an ion and then follows a stable orbit around it. The algorithm’s applicability extends to emitters with charges $Z\geq 1$. The procedure enables the correct identification of valid time-histories of the electric microfield perturbing the emitting ion, which will be used for subsequent line shape calculations. The ionization balance results obtained from the application of this algorithm are compared with an additional method based on the potential energy of the particles in the simulations, finding good agreement, therefore validating the use of this approach.
@@ I Introduction
Accurate modeling of spectral line shapes is indispensable for investigating the properties of high energy density plasmas. The analysis of emission or absorption spectra, which typically involves the comparison of experimental data with spectroscopic models, stands as a key diagnosis tool for laboratory and astrophysical plasmas. In particular, spectral line shapes, broadened and shifted by the kinetics of plasma particles and their interactions, provide relevant information about the plasma temperature and density. Spectral lines arise from atomic transitions between bound energy states of a radiating ion or neutral atom, the radiator.
@@
In the context of hot and dense plasmas the Stark effect emerges as the dominant broadening mechanism [1]. Unlike the isolated scenario, when embedded in a plasma, the radiator is subject to the electric field generated by its surrounding charged particles during the time interval in which it emits. The shape of the spectral lines in this case depends on the plasma-induced electric field, its time evolution (tied to the plasma particle dynamics) and the radiators’ ensemble.
@@
Despite significant strides in recent decades [2, 3], modeling spectral line shapes remains a formidable challenge owing to its interdisciplinary nature, requiring the integration of various ingredients from atomic physics, plasma physics, statistical mechanics, collision theory, and computational physics. Moreover, in the high-density regime, dedicated validating experiments are notably scarce, presenting a substantial obstacle to grounding theoretical models in empirical reality.
@@
Generally, two primary approaches are used for line-broadening calculations: analytic methods and computer simulations. Analytic models navigate the complexities of line broadening by incorporating assumptions and mathematical manipulations to handle the stochastic nature of line emission. One common approximation involves the separate calculation of electron and ion broadening contributions. Since ions are much heavier than electrons, they are assumed to be static over the relevant electron time scales. The analytic approach then puts the effort in obtaining an accurate representation of the electron broadening operator. In this regard, some common approximations include the simple and well-known impact approximation, or the second-order approximation as a simplified way to compute the electron broadening operator [1, 2]. Once the electron broadening contribution is computed for a given ion distribution, i.e. for a given value of the ion microfield, the line emission is integrated over the statistical distribution of the ion microfield for the conditions of interest, a facet that entails its own extensive research area in microfield probability distributions. The inherent advantages of analytic methods lie in their computational efficiency, as the time required to calculate a line shape is generally much shorter than for computer simulations, and they operate without numerical noise.
@@
On the other hand, computer simulations prove particularly advantageous in gaining insights into the line broadening problem, excelling at replicating plasma particle interactions and their impact on the state of a radiating atom or ion [4]. Simulated plasma particles traverse within a confined space, following classical trajectories in discrete timesteps. These classical particles generate a time-dependent Coulomb or Debye potential or, equivalently, a time-sequence of the electric field. The emitting atom or ion then internally evolves subjected to the effect of this electric field. By numerically solving the time-dependent Schrödinger equation, one obtains the corresponding evolution operator. It is then possible to calculate the dipolar moment of said emitter as a function of time. The Fourier transform of the autocorrelation function of this dipolar moment corresponds to the emission spectrum from the emitter under consideration. By repeating this process a sufficient number of times, the spectrum is averaged over an ensemble of emitters producing the spectral line of interest.
Simulations often simplify the problem and reduce computational time by employing straight-line path trajectories for particle movement, i.e. with the simulated particles being independent. We will refer to this type of simulations as Trivial Molecular Dynamics (TMD). TMD calculations are useful when the typical electric-interaction energies are much lower than the mean kinetic energy of particles (a weakly-coupled high temperature plasma) and, specially for non-charged emitters [5, 6, 7, 8, 9, 10].
@@
Noteworthy exceptions to this norm are Full Molecular Dynamics (FMD), in which the calculation takes into account the electric interaction of all particles in the plasma. This constitutes a more accurate (and more computationally expensive) method, which must solve Newton’s equations of motion to determine the trajectories of all particles. FMD simulations are necessary for medium and strongly-coupled plasmas and, specially when the emitter is an ion, given that its interaction with the surrounding particles in the plasma affects both its own trajectory and those of the other particles, thus modifying the time-evolution of the electric field of interest [11, 12, 13]  [^1] It is worth mentioning that there is an intermediate type of calculations, in which the perturbers are treated as independent particles, but their interaction with the emitter is taken into account explicitly. This results in the charged perturbers moving in hyperbolic trajectories when a Coulomb field is considered, or in slightly different paths if Debye screening is introduced, in which case the corresponding differential equation must be solved numerically [14]. .
While computer simulations have become a standard tool to study Stark-broadened lineshapes, there are some Physical broadening mechanisms that are still unaccounted for in these type of calculations. In this work we focus on the so-called recombination broadening, produced when a free electron within the plasma becomes bound to an emitting ion, therefore stopping its emission. While it is clear that TMD models cannot reproduce this phenomenon, as the particles are considered independent, it is possible to use FMD simulations to detect when an electron becomes bound to an ion and account for this additional broadening.
It is in this context that our work is framed. We present an algorithm to accurately track the electron capture processes within FMD computer simulations, key to obtain the correct time-sequences of the electric field within the plasma and accurately account for the additional recombination broadening.
This manuscript is structured as follows: in Section II we present the criterion for detecting electron captures within FMD simulations. An example of the application of this criterion to a computer simulation is presented in Sec. III.1, and the results are compared and validated with an additional metric in Sec. III.2. Finally, Sec. IV summarizes the conclusions and limitations of this work and explores future applications of this type of techniques.
II Criterion for detecting trapped electrons in a computer simulation
As mentioned in Section I, in FMD simulations, the particle trajectories are solved by accounting for the Coulombian interaction between all particles in the plasma. As the simulation evolves, it becomes possible that the binding energy (negative) between an electron and an ion becomes too large - in absolute value - with respect to their kinetic energies (positive), so that the electron becomes trapped by the ion. The two particles in this pair can therefore no longer be considered free and must be labeled as a neutral atom, or as an ion with a positive charge lower than that of the free ion.
@@ II Criterion for detecting trapped electrons in a computer simulation
This circumstance was extensively discussed in our previous work [15], where we developed a statistical model for the equilibrium of the simulated plasma, taking into account the process of electron capture and release by the ions (recombination and ionization respectively). For this model, it was necessary to establish a clear criterion to determine when an electron-ion pair must be considered bound, and when those particles must be considered free. While in Gigosos et al. [15] ions of charge +e+e were used (where $e$ is the electron charge), in this work we are interested in plasmas with charge +Ze+Ze, with $Z>1$. The purpose of this work is then to extend this criterion to situations in which the number of trapped electrons can exceed $1$.
@@
Furthermore, there is another important reason that requires knowing when an ion has captured an electron and when it can be considered free. Let us consider a FMD simulation in which we are interested in calculating the spectrum from an ionic emitter with charge +Ze+Ze. To that end, during the simulation process, we must register the temporal sequence of the electric field seen by that charged emitter, which is produced by the ions and free electrons surrounding it. If, at a given moment, an electron is captured by the ion of consideration, from that moment on, the electron will orbit around it at a short distance. Therefore, the electric field felt by that emitter will be an oscillating function with a large amplitude (since the oscillating electron is very close), which spoils the field sequence.
In addition, the capturing ion, originally with charge +Ze+Ze becomes now, along with the trapped electron, an ion of charge +(Z1)e+(Z-1)e and can therefore no longer be considered a contributor to the spectral line in which we are interested. The moment the electron is captured determines when the ion changes its species and therefore the emission coherence is abruptly lost. This causes a broadening of the line by the recombination process, which highlights one the advantages of these calculations [13]. As a consequence, in practice, the corresponding time-sequence of the electric field is cut in the simulation process. It is then necessary to wait for the corresponding ion to recover its original charge in order to continue using it for the calculation.
@@
To correctly select the valid field sequences we choose a global criterion, in which whether an ion has trapped an electron or not, not only depends on its conditions at a given time, but also on its history. For a given ion, we store (for each step of the simulation) its potential energy, the electric field at its position and the label (an index identifying each electron in the simulation) of the $Z+1$ closest electrons, where $Ze$ is the charge of the emitter.
In order to correctly calculate the potential energy, since the interaction between particles is Coulombian, it is necessary to apply a regularization to the interaction potential to avoid divergencies at short distances. While several procedures for this regularization have been reported (see Bonitz et al. [18] and references therein), in this work we will use that described in Gigosos et al. [15], which considers a parabolic potential for interparticle distances below a given value $a$ and hyperbolic for larger distances. This method naturally accounts for the diffraction of electrons around an ion and determines an effective atom size. We used the following expression for $Z>1$ ions, rescaled from Gigosos et al. [15]
$V_{ie}(r)=\begin{cases}V_{i}\left[\frac{1}{3}\left(\frac{r}{a}\right)^{2}-1\right],&\quad r\leq a\\ -\frac{Ze^{2}}{4\pi\varepsilon_{0}r},&\quad a<r\leq R_{I}\\ 0,&\quad r>R_{I}\end{cases}$ (1)
where $V_{i}$ is the ionization energy corresponding to the ion of charge $Z$ and $R_{I}$ is the radius of the interaction sphere, defined as half the size of the cubic simulation box. For ions with $Z>1$, $a$ is a parameter given by
$a=\frac{3}{2}\cdot\frac{Ze^{2}}{4\pi\varepsilon_{0}V_{i}}.$ (2)
The first step then is to track down the moments when the potential energy of the emitter is lower than a given threshold energy $\mathscr{E}_{th}$. In particular, to keep consistency with the original criterion for $Z=1$ atoms, we choose
$\mathscr{E}_{th}=-\frac{2}{3}V_{i},$ (3)
which corresponds to the potential energy of an ion with an electron at a distance $a$. We note here that the exact value of $\mathscr{E}_{th}$, while important, is not critical for this criterion, as we will show in Section III.1.
Whenever the ion energy is lower than $\mathscr{E}_{th}$, the electron that is closest to it (its first neighbor) is identified. Then, the time interval during which that electron has been among the $Z+1$ closest neighbors is considered. If this interval is longer than a given threshold value $\tau_{\text{bound}}$, then the average total energy of the electron-ion pair in that interval is calculated. If the pair has a negative total energy during that interval, then the electron is considered to be bound for the whole interval. If not, it is considered that the negative energy seen by the emitter is caused by a strong collision, since the electron causing it did not remain around the ion. It is worth noting that the $Z+1$ closest neighbors are considered, since it is possible that a fast electron gets closer to the ion core than other bound electrons for a short period of time, without freeing them.
Since the full time-history of the ion is needed in order to check whether an electron is bound or free, its history must be saved so that it can be accessed when the energy criterion is met, and the correct field sequences can be extracted.
Determining the timescale for bound electrons
@@
$\tau_{T}=\frac{2\pi r_{T}}{v_{T}}=\frac{Ze^{2}}{2\varepsilon_{0}}\cdot\frac{m_{e}^{1/2}}{(2k_{B}T_{e})^{3/2}}.$ (6)
This expression can be simplified by defining the typical timescale of electron dynamics $t_{0}$ as the time it takes an electron moving with the characteristic thermal velocity $v_{T}$ to travel the typical distance between electrons ($r_{e}=(4\pi n_{e}/3)^{-1/3}$). Using this definition, one gets
$t_{0}=\frac{r_{e}}{v_{T}}=\left(\frac{3}{4\pi n_{e}}\right)^{1/3}\cdot\left(\frac{m_{e}}{2k_{B}T_{e}}\right)^{1/2},$ (7)
which yields
$\tau_{T}=\frac{\pi}{3}Z\rho^{2}t_{0},$ (8)
where $\rho$ is the coupling degree of the plasma, given by the relation between the typical distance between electrons and the Debye radius.
We take this characteristic orbit time as a basis to consider whether an electron is bound to an ion. Although one might be tempted to consider $\tau_{T}$ as the temporal threshold - that is, if an electron is within the $Z+1$ closest neighbors of a given ion for at least $\tau_{T}$, and its total energy is negative, it is considered bound - this criterion would overestimate the number of bound electrons, since an electron can loop around an ion without being bound. To account for this possibility, we take $\tau_{\text{bound}}=3\tau_{T}$ as the threshold for considering an electron as bound to a given ion.
We note here that the above expression for $\tau_{T}$ was obtained for a hyperbolic potential. However, as indicated in Eqn. 1, in our simulations, the potential becomes parabolic for distances $r<a$, and thus, for this expression to be valid, it is necessary that $r_{T}>a$. By comparing the above expressions it can easily be seen that this condition is equivalent to
@@
Figure 1: Time-history of a plasma with $\rho=0.6$ and emitters with charge 2. From top to bottom: potential energy of an emitter as a function of time, indices of the three closest electrons at each moment, and electric field. The shaded regions correspond to the moments when a particle is trapped. The potential energy is given in units of $\mathscr{E}_{0}$, a characteristic energy of the simulation, and the electric field in units of $E_{0}$, the electric field produced by a single electron at the typical distance $r_{e}$.
III Results and discussion
III.1 Application of the criterion